rvc (rvc):

if one of the line represented by the equation ax^2+2hxy+by^2=0 is coincident with one of the line represented by a'x^2+2h'xy+b'y^2=0 then

OpenStudy (anonymous):

same partial derivatives..

rvc (rvc):

@SolomonZelman here it is

OpenStudy (perl):

these equations look like quadratics, but if you want to get a line, they will need to be perfect squares

rvc (rvc):

(ab'-a'b)^2=4(ah'-a'h)(hb'-h'b) (ab'+a'b)^2=4(ah'-a'h)(hb'-h'b) (ab'-a'b)^2=(ah'-a'h)(hb'-h'b) (a'b'-ab)^2=(ah-a'h')(hb-h'b') these are the options

rvc (rvc):

@rational @Jhannybean @hartnn @perl help

rvc (rvc):

@TheSmartOne @Kainui @iambatman help me please

rvc (rvc):

@dan815

rvc (rvc):

@sammixboo

sammixboo (sammixboo):

Ah sorry can't help ;-;

OpenStudy (domebotnos):

@perl @SolomonZelman

OpenStudy (anonymous):

It's 0046, local time. Have to bail out now. Sorry.

rvc (rvc):

@ParthKohli @StudyGurl14 @EclipsedStar @rational

rvc (rvc):

@dan815 may u help

OpenStudy (dan815):



OpenStudy (rational):

did you make it with a script ? XD

OpenStudy (dan815):

: )

OpenStudy (rational):

send me the script

OpenStudy (rational):

mouse listener or whatever..

rvc (rvc):

hmmm

rvc (rvc):

@satellite73 im sure you can help me

rvc (rvc):

@satellite73 need help

rvc (rvc):

@uri

rvc (rvc):

@Vincent-Lyon.Fr

rvc (rvc):

@divu.mkr

rvc (rvc):

@bibby @King.Void.

rvc (rvc):

@ikram002p @ganeshie8

rvc (rvc):

@campbell_st @Callisto

rvc (rvc):

@Somy have a look :)

OpenStudy (anonymous):

partial derivate will give you two lines...out of four lines a pair will be useless rest two will be useful. solve them as system of linear equations using cramer rule you could get the idea by checking the options..they simply suggest of cramer rule

rvc (rvc):

ha can u solve this question pls @divu.mkr

rvc (rvc):

@YanaSidlinskiy

rvc (rvc):

@ikram002p plssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssssss

rvc (rvc):

@ashwinram

rvc (rvc):

hmm

OpenStudy (mimi_x3):

if one of the line represented by the equation ax^2+2hxy+by^2=0 is coincident with one of the line represented by a'x^2+2h'xy+b'y^2=0 then

OpenStudy (mimi_x3):

sharing lines

rvc (rvc):

no the options are mentioned

OpenStudy (mimi_x3):

that is the definition of coincidence

OpenStudy (mimi_x3):

lines on top of each other

OpenStudy (rational):

\[(y+m_1x)(y+m_2x) = 0\tag{1}\] \[(y+m_1x)(y+m_3x) = 0\tag{2}\]

OpenStudy (mimi_x3):

if its a perfect square u can get a y=mx relationship

rvc (rvc):

hey wait

rvc (rvc):

im posting this question again in another section

rvc (rvc):

closing this question

OpenStudy (rational):

we can always break ax^2+bxy+by^2 = 0 as (y+m1x)(y+m2x) = 0 as they represent a pair of straight lines passing through origin